OpenStudy (anonymous):

a rectangle has a diagonal of 10 and length of 8 what is the with?

OpenStudy (anonymous):

if you draw it , you'll notice that you have a right triangle, so use the pythagorous theorem:\[c^2 = x^2 + y^2 \] but in this case we have C, so you have to find y : \[y^2 = c^2 - x^2 =\sqrt{10^2 - 8^2}= \sqrt{164}= 2\sqrt{41}\] Correct me if I'm wrong ^_^

OpenStudy (anonymous):

w = l = 2sqrt(41)

OpenStudy (anonymous):

no your right it looks right to me now that i look at it

OpenStudy (anonymous):

lol, okay then :)